Randomized Algorithms - Course Review

K. Subramani LCSEE, West Virginia University, Morgantown, WV {ksmani@csee.wvu.edu}

March 6, 2012

1. Does the instructor address your concerns in timely fashion?

(a) Yes.

(b) No.

(c) Most of the time.

2. Is the instructor helpful?

(a) Yes.

(b) No.

(c) Most of the time.

(d) Sometimes.

3. Have you sought help?

(a) Yes.

(b) No.

4. How much time do you spend on this subject outside of class?

(a) None.

(b) 1-2 hours per week.

(c) 4-5 hours per week.

5. Are you learning new material in this course? (a) A lot.

(b) Some.

(c) No.

6. Are your critical thinking skills improving through this course?

(a) Yes, significantly.

(b) Somewhat.

(c) No.

7. Have you missed more than three classes?

(a) Yes.

(b) No.

8. Do you carefully review the material covered in previous classes before attending class?

(a) Never.

(b) Rarely.

(c) Most of the time.

(d) Always.

9. How would you characterize the content of the course?

(a) Easy.

(b) Difficult.

(c) Difficult but rigorous.

10. How would you rate this instructor vis-a-vis other instructors that you have had?

(a) Better than other instructors.

(b) As good as other instructors.

(c) Worse than other instructors.

11. How would you rate this course vis-a-vis other courses that you have taken?

(a) Better than other courses.

(b) As good as other courses.

(c) Worse than other courses.

12. Are the homework problems helping you appreciate the course?

(a) Yes.

(b) No.

(c) For the most part.

13. Are the homework problems fair?

(a) Yes.

(b) No.

(c) For the most part.

14. Is the grading fair?

(a) Yes.

(b) No.

(c) For the most part.

15. Are you getting timely feedback on your work?

(a) Yes.

(b) No.

(c) For the most part.

16. Do you believe that the non-uniformity of student backgrounds impacts the instructor’s effectiveness?

(a) Yes.

(b) No.

17. Briefly explain what your notion of an effective instructor is:

18. Does the instructor meet your notion of an effective instructor:

(a) Always.

(b) Never.

(c) Most of the time.

(d) Sometimes.

19. What aspects of the course would you like to change and in what way? (More detail the better!)